OpenStudy (aaronandyson):

Find the equation of a line hose y-intercept and slope have the same numerical value and pass through (4,-5)

OpenStudy (aaronandyson):

*whose

OpenStudy (tootzrll):

first, lets start with a graph: 

OpenStudy (tootzrll):

now that we have a graph, lets simplify the problem. what we have to do is find the point on which certain coordinates intersect, then plot the course. sounds complicated, maybe.

OpenStudy (kropot72):

The general equation for a line is: y = mx + b where m is the slope and b is the y-intercept. In this case we can write: y = mx + m ............(1) Plugging the given values of the coordinates of the point into (1) we get: -5 = 4m + m ..........(2) If you solve equation (2) to find the value of m, that value can be plugged into equation (1) to give the equation of the line.

OpenStudy (aaronandyson):

x+y+1 =0?

OpenStudy (kropot72):

@AaronAndyson What is your value for m?

OpenStudy (aaronandyson):

-1

OpenStudy (kropot72):

Correct. So plug that value into equation (2).

OpenStudy (kropot72):

y = -x - 1 is the required equation.